# Bayesian Optimisation of Breit-Wheeler Pair Production in Simulated Laser Experiments

### v1

@@ Abstract
High laser intensities enable the production of electron-positron pairs from bright gamma rays passing through strong fields. Potentially the most promising approach for all-optical experiments in the near term uses dense but higher divergence electron beams from laser wakefield acceleration to produce gamma rays through inverse Compton scattering. Achieving many-photon collisions between these gamma rays and the high intensity laser pulse in practice is extremely difficult, however, due to significant shot-to-shot jitter in laser pointing and timing.
We model these practical difficulties using simulated Monte-Carlo experiments. By using a more efficient algorithm for sampling infrequent pair production with particle splitting, we enable the exploration of a multi-dimensional parameter space. Using Gaussian Process Regression we then efficiently find optimal conditions for maximising pair production by changing the laser spot size, the energy in the colliding beam, and the stand-off distance between the laser wakefield accelerator and the focus of the colliding laser pulse. We find that the optimal stand-off distance increases with the degree of laser jitter and that the best conditions for producing electron-positron pairs are not the same as the best conditions for maximising the energy in the gamma rays. With 100 J of laser energy, we estimate rates of pair production of around 1 pair per 100 electrons are achievable even with jitter of 10s of microns and 10s of femtoseconds.
@@ 1 Introduction
Laser facilities continue to push the intensity frontier to new records, in recent years achieving peak powers of $10\,\mathrm{PW}$[1] and peak intensities of $10^{23}\,\mathrm{Wcm^{-2}}$[2]. Several studies have explored the possibility of creating matter from pure light in sufficiently strong laser fields[3, 4, 5]. In the centre-of-mass frame of this many-photon interaction, the field strength approaches the Schwinger field [6], enabling the production of position-electron pairs in vacuum through the non-linear Breit-Wheeler process [7]. Previous experiments have been able to reach this regime using a $47\,\mathrm{GeV}$ electron beam from SLAC colliding with an $0.4\,\mathrm{TW}$ laser pulse to produce around $100$ positron-electron pairs from around $10^{10}$ electrons [8]. This first produces bright $\gamma$-rays from inverse Compton scattering [9], before the $\gamma$-rays themselves interact with the laser field. Another approach has been to use the strong atomic fields present within periodic crystal lattices [10, 11], with a $200\,\mathrm{GeV}$ electron beam driving trident pair production in a germanium crystal. Now, several laser facilities aim to utilise unprecedented laser powers and produce positron-electron pairs in vacuum by using laser-driven plasma wakefield accelerators[12, 13].
@@
Here, we find optimum parameters for electron-positron pair production in all-optical experiments by employing two techniques for performing many simulations with higher computational efficiency. First, we describe a particle splitting algorithm and show how it requires changes to the way we model pair production. This accurately calculates the rate of pair production for rare events without significantly increasing the runtime, allowing us to rapidly conduct many simulated experiments, including the effect of jitter. Second, we use a Bayesian optimisation scheme built on Gaussian process regression to efficiently explore the parameter space and find the maximum rate of pair production. This scheme is significantly faster than a brute force grid search and builds a fast surrogate model for estimating the rate of pair production across the domain. We show that more position-electron pairs can be produced by combining the tightest possible laser focus with a few-cm-scale stand-off distance between the laser wakefield accelerator and the interaction point. This ameliorates the effect of jitter while retaining the high field strength required for pair production.
@@ 2.1 Pair Production Algorithm
In both Monte-Carlo and Particle-In-Cell codes, the emission of high energy photons and the production of positron-electron pairs are treated as discrete random and independent events which can occur in any fixed time interval. As such, they can be modelled using a Poisson distribution, where the average number of emissions per particle in a given timestep $\delta t$ are given by:
$\lambda_{\gamma}=\frac{dN_{\gamma}(\chi_{e})}{dt}\delta t\\ \lambda_{e^{+},e^{-}}=\frac{dN_{e^{+},e^{-}}(\chi_{\gamma})}{dt}\delta t,$ (1)
@@
Often, in order to reduce the computational expense of sampling from a Poisson distribution at every timestep, codes use an optical depth approach. If the emission rate is sufficiently low, the chances of two or more emissions in a single timestep is negligible and the probability of a single particle being emitted on a timestep before $t_{N}$ becomes $P(t_{0}<t<t_{N})\approx 1-\exp\left[-\sum^{N}_{n=0}\lambda(t_{n})\right]$. Now a uniformly distributed random number in the range $0\leq r<1$ is chosen to represent $P(t)$ and an emission occurs on the first timestep when $1-\exp\left[-\sum_{n}\lambda(t_{n})\right]\leq r$, which is when the accumulated optical depth reaches $\sum_{n}\lambda(t_{n})\geq-\ln(1-r)$. Only one random number must be chosen for each emission event, calculated the moment a particle is created or the moment after an emission takes place, and a single running total tracks the optical depth. The optical depth method therefore efficiently models rare emissions which occur singly and consecutively.
@@ 2.2 Particle Splitting Algorithms
We therefore propose introducing particle splitting into the calculation of Breit-Wheeler pair production in Monte-Carlo and PIC codes. In PIC codes, each macroparticle represents a large number of underlying real particles, which move around the simulation grid together and sample the underlying distribution in $(\mathbf{x},\mathbf{p})$ space with a given weight $w$. In a similar way, each macroparticle will move around as before, but will independently sample the chances of producing a positron-electron pair many times per timestep at a splitting number $u$, making the average number of emissions per timestep now equal to:
@@
The weight of any pairs produced will be reduced proportionately, from $w$ to $1/u$, and the weight of the photon will be slightly reduced to account for the decay, from $w$ to w1/uw-1/u. If this weight falls to zero the photon is removed from the simulation. Particle splitting does not affect the number of photon particles and does not significantly increase the cost of the particle push, only affecting the number of pair particles. In this way, emission events can be made more common without greatly increasing the runtime, more efficiently modelling pair production.
@@
However, there is a risk to naively using the optical depth algorithm with an increased emission rate for this purpose. Firstly, if the degree of particle splitting $u$ is sufficiently high that $\lambda_{e^{+},e^{-}}\sim 1$, there is now a non-negligible chance that more than one pair will be produced per timestep, giving $P(0<t<t_{N})\neq 1-\exp\left[-\sum^{N}_{n=1}\lambda(t_{n})\right]$. Secondly, even when this assumption does hold, the optical depth is refreshed after each emission using a new random number, wiping the history of the particle and making emissions no longer independent. Using particle splitting while using a single optical depth per particle can therefore give a drastically different estimate for the rate of pair production.
@@
Instead, we propose sampling the full Poisson distribution on every timestep to model particle splitting. The average number of emissions in each timestep is still increased by a factor $u$ but now any number of emissions are possible in any timestep. If at least one emission occurs in a single timestep, $n>0$, a single pair of electron and positron macroparticles can be produced, each with a weight $n/u$. Alternatively, if a larger number of pair particles are desired, for instance to better sample an energy distribution, $n$ macroparticles can be produced, each with a weight $1/u$. In either case, the weight of the photon is reduced to wn/uw-n/u. This ensures that pair production events are independent while increasing the efficiency of sampling the probability of pair production.
@@
Another approach is to add the new optical depth onto the previous value, including the ‘overshoot’ by which the optical depth fell below zero. Another emission occurs when this additive optical depth falls below 0. As the optical depth becomes the sum of all the random numbers $\sum_{i}\ln(1-r_{i})$, minus the total accumulated event path length $\sum_{n}\lambda t_{n}$, this maintains both the history of the particle and the independence of the emissions, with a single unique random number for each emission event. In this case, the particle weight is not updated after each emission. Instead, the particle is killed with a probability of $1/u$, such that at higher upscalings the majority of particles are unaffected by emissions. On the ensemble average, however, this reduces the population of photons to account for energy conservation.
@@ 2.3 Particle Splitting Simulations
We tested the three techniques in Monte-Carlo simulations of a head-on collision between high energy photons and an intense laser pulse using the QED-Cascade code [14]. A $1\,\mathrm{PW}$ pulse with $25\,\mathrm{J}$ in $25\,\mathrm{fs}$ was simulated, focussed to a spot waist of $w_{0}=2\,\mathrm{\mu m}$ and a peak intensity of $2.2\times 10^{22}\,\mathrm{Wcm^{-2}}$ or $a_{0}\approx 100$. $1000$ photon particles, each with an initial weight of $w=1$, were collided with the laser pulse at an angle of $15^{\circ}$, intersecting the peak laser intensity at best focus. The photon beam was a mono-energetic and divergence-free point source with energies varying from $100\,\mathrm{MeV}$ up to $1\,\mathrm{GeV}$. 10 simulations were run at each photon energy, with the results shown in Fig. 1. This shows both the number of pairs produced and the change in the predicted energy spectrum of the positrons.
At photon energies below $300\,\mathrm{MeV}$, the rate of pair production is sufficiently small (below $10^{-4}$ positrons per electron) that the original simulations without splitting do not produce any positron-electron pairs, leaving the simulations unable to capture the Breit-Wheeler process. Even above this, very small numbers of pairs are produced, giving significant variation from shot noise, as shown by large error bars at each point. At $400\,\mathrm{MeV}$, only four positron macroparticles were produced, giving large shot noise and making it hard to resolve the positron energy distribution.
@@
Using both of the corrected splitting methods, returning to sampling the Poisson distribution and adding new optical depths onto the cumulative total, the simulations are able to accurately resolve the pair production rate both at photon energies as low as $150\,\mathrm{MeV}$, and up to $1\,\mathrm{GeV}$. Here, the shot noise is lower and the larger number of macroparticles (100,000 at $1\,\mathrm{GeV}$) enables fine resolution of the pair production rates and the positron energy distributions. Both corrected splitting methods are therefore effective for simulating Breit-Wheeler pair production across a wider range of photon energies, even with a relatively small number of photon macroparticles. For the rest of the paper, we focus on the Poisson approach, which is closer to the true physical picture at the cost of a higher runtime.
Figure 1: a) Rate of electron-positron pair production in a tightly focussed 1 PW beam against photon energy, comparing results without splitting (blue circles), with naive splitting (red xs), with Poisson-sampling splitting (purple crosses), and with additive optical depths (green diamonds). b) Positron spectra predicted without splitting (blue) and with naive (red) and corrected splitting (purple), for photons at 400 MeV (i) and 1 GeV (ii).

### v2

# Bayesian Optimisation of Non-linear Breit-Wheeler Pair Production in Simulated Laser Experiments

@@ Abstract
High laser intensities enable the production of electron-positron pairs from bright gamma rays passing through strong fields. Potentially the most promising approach for all-optical experiments in the near term uses dense but higher divergence electron beams from laser wakefield acceleration to produce gamma rays through inverse Compton scattering. Achieving many-photon collisions between these gamma rays and the high intensity laser pulse in practice is extremely difficult, however, due to significant shot-to-shot jitter in laser pointing and timing. In practice, this jitter reduces the yield of electron-positron pairs by orders of magnitude.
We model these practical difficulties using simulated Monte-Carlo experiments. By using a more efficient algorithm for sampling infrequent pair production with particle splitting, we enable the exploration of a multi-dimensional parameter space. Using Gaussian Process Regression we then efficiently find optimal conditions for maximising pair production by changing the laser spot size, the energy in the colliding beam, and the stand-off distance between the laser wakefield accelerator and the focus of the colliding laser pulse. We find that the optimal stand-off distance increases with the degree of laser jitter and that the best conditions for producing electron-positron pairs are not the same as the best conditions for maximising the energy in the gamma rays. With 100 J of laser energy, we estimate realistic rates of pair production of around 1 pair per 100 electrons are achievable even with jitter of 10s of microns and 10s of femtoseconds.
@@ 1 Introduction
Laser facilities continue to push the intensity frontier to new records, in recent years achieving peak powers of $10\,\mathrm{PW}$[1] and peak intensities of $10^{23}\,\mathrm{Wcm^{-2}}$[2]. Several studies have explored the possibility of creating matter from pure light in sufficiently strong laser fields[3, 4, 5]. In the centre-of-mass frame of this many-photon interaction, the field strength approaches the Schwinger field [6], enabling the production of positron-electron pairs in vacuum through the non-linear Breit-Wheeler process [7]. Previous experiments have been able to reach this regime using a $47\,\mathrm{GeV}$ electron beam from SLAC colliding with an $0.4\,\mathrm{TW}$ laser pulse to produce around $100$ positron-electron pairs from around $10^{10}$ electrons [8]. This first produces bright $\gamma$-rays from inverse Compton scattering [9], before the $\gamma$-rays themselves interact with the laser field. Another approach has been to use the strong atomic fields present within periodic crystal lattices [10, 11], with a $200\,\mathrm{GeV}$ electron beam driving trident pair production in a germanium crystal. Now, several laser facilities aim to utilise unprecedented laser powers and produce positron-electron pairs in vacuum by using laser-driven plasma wakefield accelerators[12, 13].
@@
Several Monte-Carlo and Particle-In-Cell codes have already increased computational efficiency by introducing particle merging [14, 15, 16] and splitting [17, 18, 19, 20] algorithms. Merging techniques are useful when lots of particles are being produced, using too much computational memory, whereas splitting techniques are useful when few particles are produced, allowing codes to explore sparse areas of the parameter space and reduce the variance in the results. With Breit-Wheeler pair production a rare process with current laser parameters, particle splitting is a promising approach. However, most particle splitting algorithms such as Refs. [18, 20] split the primary particle (in our case photons), which we generally have many of already, unnecessarily increasing computation in the particle push or tracking. On the other hand, algorithms like Refs. [17, 19] split the secondary particle (in our case a positron-electron pair) only after it has been produced, which is unhelpful if the chances of it being produced in the first place are so small. We therefore want to perform particle splitting during the pair production step, which is a novel and computationally efficient approach but risks affecting the algorithm’s accuracy.
Here, we find optimum parameters for electron-positron pair production in all-optical experiments by employing two techniques for performing many simulations with higher computational efficiency. First, we describe a particle splitting algorithm that performs the split during pair production, and show how it requires changes to the way we model pair production. This accurately calculates the rate of pair production for rare events without significantly increasing the runtime, allowing us to rapidly conduct many simulated experiments, including the effect of jitter. Second, we use a Bayesian optimisation scheme built on Gaussian process regression to efficiently explore the parameter space and find the maximum rate of pair production. This scheme is significantly faster than a brute force grid search and builds a fast surrogate model for estimating the rate of pair production across the domain. We show that more positron-electron pairs can be produced by combining the tightest possible laser focus with a few-cm-scale stand-off distance between the laser wakefield accelerator and the interaction point. This ameliorates the effect of jitter while retaining the high field strength required for pair production.
@@ 2.1 Pair Production Algorithm
In both Monte-Carlo and Particle-In-Cell (PIC) codes, the emission of high energy photons and the production of positron-electron pairs are treated as discrete random and independent events which can occur in any fixed time interval. As such, they can be modelled using a Poisson distribution, where the average number of emissions per particle in a given timestep $\delta t$ are given by:
$\lambda_{\gamma}$ $=\frac{dN_{\gamma}(\chi_{e})}{dt}\delta t$ (1)
$\lambda_{e^{+},e^{-}}$ $=\frac{dN_{e^{+},e^{-}}(\chi_{\gamma})}{dt}\delta t,$ (2)
@@
Often, in order to reduce the computational expense of sampling from a Poisson distribution at every timestep, codes use an optical depth approach (for instance in the EPOCH PIC code [21, 22]). If the emission rate is sufficiently low, the chances of two or more emissions in a single timestep is negligible and the probability of a single particle being emitted on a timestep before $t_{N}$ becomes:
$P_{\mathrm{emission}}(t_{0}<t<t_{N})$ $\approx 1-P_{\mathrm{no~emission}}(t_{0}<t<t_{N})$
$\approx 1-\exp\left[-\sum^{N}_{n=0}\lambda(t_{n})\right].$
Now a uniformly distributed random number in the range $0\leq r<1$ is chosen to represent $P(t)$ and an emission occurs on the first timestep when $1-\exp\left[-\sum_{n}\lambda(t_{n})\right]\leq r$, which is when the accumulated optical depth reaches $\sum_{n}\lambda(t_{n})\geq-\ln(1-r)$. Only one random number must be chosen for each emission event, calculated the moment a particle is created or the moment after an emission takes place, and a single running total tracks the optical depth. The optical depth method therefore efficiently models rare emissions which occur singly and consecutively.
@@ 2.2 Particle Splitting Algorithms
We therefore want to introduce particle splitting into the calculation of Breit-Wheeler pair production in Monte-Carlo and PIC codes, during the pair production step rather than before or afterwards. In PIC codes, each macroparticle represents a large number of underlying real particles, which move around the simulation grid together and sample the underlying distribution in $(\mathbf{x},\mathbf{p})$ space with a given weight $w$. When PIC codes model random emission processes they typically use the method discussed in Section 2.1, sampling the probability distribution function once per timestep and creating child particles which also each have a weight of $w$. Here we describe how we can let each macroparticle move around as before, but instead independently sample the chances of producing a positron-electron pair many times per timestep at a splitting number $uw$, making the average number of emissions per timestep now equal to:
@@
The weight of any pairs produced will be reduced proportionately, from $w$ to $1/u$, and the weight of the photon will be slightly reduced to account for the decay, from $w$ to w1/uw-1/u. If this weight falls to zero the photon is removed from the simulation. Particle splitting does not affect the number of parent particles and does not significantly increase the cost of the particle push, only affecting the number of pair particles. In this way, emission events can be made more common without greatly increasing the runtime, more efficiently modelling pair production.
@@
However, there is a risk to naively using the optical depth algorithm with a high rate of particle splitting for this purpose. Firstly, if the degree of particle splitting $u$ is sufficiently high that $\lambda_{e^{+},e^{-}}\sim 1$, there is now a non-negligible chance that more than one pair will be produced per timestep, giving $P_{\mathrm{emission}}(0<t<t_{N})\neq 1-\exp\left[-\sum^{N}_{n=1}\lambda(t_{n})\right]$. Secondly, even when only one or fewer pairs is produced in a timestep, the optical depth is refreshed after each emission using a new random number. This wipes the history for all particles represented by the macroparticle, not just the fraction $1/u$ that has been emitted, making emissions no longer independent. Without particle splitting, neither of these are issues, as photons are killed immediately upon producing just one electron-positron pair. Using particle splitting while using a single optical depth per particle, however, can give a drastically different estimate for the rate of pair production.
This problem can be resolved in three different ways, with different degrees of accuracy:
@@
I - Sampling the full Poisson distribution on every timestep to model particle splitting. The average number of emissions in each timestep is still increased by a factor $u$ but now any number of emissions are possible in any timestep. If at least one emission occurs in a single timestep, $n>0$, a single pair of electron and positron macroparticles can be produced, each with a weight $n/u$. Alternatively, if a larger number of pair particles are desired, for instance to better sample an energy distribution, $n$ macroparticles can be produced, each with a weight $1/u$. In either case, the weight of the photon is reduced to wn/uw-n/u. This is a relatively simple extension for codes like PTARMIGAN [23], which already calculate the emission probability on every timestep. It ensures that pair production events are independent and conserves energy on every emission, while allowing for an arbitrarily large degree of particle splitting.
@@
II - Adding the new random optical depth onto the previous value, including the ‘overshoot’ by which the optical depth fell below zero. Another emission occurs when this additive optical depth falls below 0 on a subsequent timestep. As the optical depth becomes the sum of all the random numbers $\sum_{i}\ln(1-r_{i})$, minus the total accumulated event path length $\sum_{n}\lambda t_{n}$, this maintains both the history of the particle and the independence of the emissions, while using only a single random number for each emission event. In this case, the parent particle weight is not updated after each emission, in order to make sure the emissions are still independent. Instead, the particle is killed with a probability of $1/u$, such that at higher splitting rates the majority of particles are unaffected by emissions. This is a natural extension for codes that use optical depths, like EPOCH [22] or Smilei [24]. It makes pair production events independent again while conserving energy on average.
III - ‘Sub-cycling’ the addition of multiple random optical depths onto the previous value to represent multiple emissions. Typically the ‘additive optical depth’ approach will not allow for more than one emission per timestep, so will also fail for high rates of pair production $\lambda\sim 1$, where more than one pair can be produced by a single macroparticle per timestep. However, it’s possible to repeat the process multiple times, continuing to compute emissions so long as the optical depth remains below zero. In practice, this repeated optical depth approach is the technique used by many Poisson sampling algorithms when the mean is not much greater than one, and so if $\lambda$ is around one or less this sub-cyling approach is still more efficient than Poisson sampling. However, sub-cycling still does not update the parent particle weight after each emission and so only conserves energy on the ensemble average.
The four approaches are summarised in Table 1. The most expensive and most accurate approach is full Poisson sampling, which allows for multiple emissions per timestep at very high field strengths and conserves energy on every emission, but requires calculating a random number every timestep (more than one random number if there is more than one emission) and is therefore more expensive. It also requires modelling particles with a wide range of different weights (wn/uw-n/u and $n/u$ for any integer $n$), which can add computational challenges. The naive optical depth approach and additive optical depth approaches are equally computationally cheap, but using an additive optical depth extends the range where upscaling is valid to areas where a particle will have more than one emission over the course of the simulation, at the cost of only conserving energy on average. Subcycling extends the range of applicability still further by allowing multiple emissions per timestep, at the cost of more computation only when multiple emissions per timestep are required. From here on we will only use the version of the additive optical depth with sub-cycling included. All of these optical depth approaches have simpler particle weights, either ‘big’ macroparticles with weight $w$ or ‘small’ macroparticles with weight $1/u$. This makes subsequent numerical behaviour simpler, such as by ignoring ‘small’ particles in subsequent collisions.
Algorithm Computation Valid when: Simple Weights? Conserves energy?
Naive optical depth Min($\lambda$,11) $\lambda\ll 1$ Yes Exactly
Additive optical depth Min($\lambda$,11) $\lambda\lessapprox 1$ Yes On average
Additive plus subcycling $\lambda$ Any $\lambda$ Yes On average
Poisson Max($\lambda$,11) Any $\lambda$ No Exactly
Table 1: Summary of the different particle splitting techniques for an average number of pairs produced per timestep $\lambda=uw\frac{dN_{+}}{dt}\delta t$, including a particle splitting rate $u$. Columns describe the approximate computational cost (as the average number of random numbers required per macroparticle per timestep), the range of $\lambda$ where the algorithm is appropriate, whether the algorithm keeps weights simple (either $w$ or $1/u$), and the energy conservation of the algorithm.
@@ 2.3 Particle Splitting Simulations
We tested the three techniques for particle splitting in Monte-Carlo simulations of a head-on collision between high energy photons and an intense laser pulse using the QED-Cascade code [25, 26]. A $1\,\mathrm{PW}$ pulse with $25\,\mathrm{J}$ in $25\,\mathrm{fs}$ was simulated, focussed to a spot waist of $w_{0}=2\,\mathrm{\mu m}$ and a peak intensity of $2.2\times 10^{22}\,\mathrm{Wcm^{-2}}$ or $a_{0}\approx 100$. $1000$ photon particles, each with an initial weight of $w=1$, were collided with the laser pulse at an angle of $15^{\circ}$, intersecting the peak laser intensity at best focus. The photon beam was a mono-energetic and divergence-free point source with energies varying from $100\,\mathrm{MeV}$ up to $1\,\mathrm{GeV}$. 10 simulations were run at each photon energy, with the results shown in Fig. 1. This shows both the number of pairs produced and the change in the predicted energy spectrum of the positrons.
At photon energies below $300\,\mathrm{MeV}$, the rate of pair production is sufficiently small (below $10^{-4}$ positrons per photon) that the original simulations without splitting do not produce any positron-electron pairs, leaving the simulations unable to capture the Breit-Wheeler process. Even above this, very small numbers of pairs are produced, giving significant variation from shot noise, as shown by large error bars at each point. At $400\,\mathrm{MeV}$, only four positron macroparticles were produced, giving large shot noise and making it hard to resolve the positron energy distribution. For convergence testing at these low energies we require simulations with more photon macroparticles, described later (see Fig. 2).
@@
Using both of the corrected splitting methods, returning to sampling the Poisson distribution and adding new optical depths onto the cumulative total, the simulations are able to accurately resolve the pair production rate both at photon energies as low as $150\,\mathrm{MeV}$, and up to $1\,\mathrm{GeV}$. Here, the shot noise is lower and the larger number of positron macroparticles (100,000 at $1\,\mathrm{GeV}$) enables fine resolution of the pair production rates and the positron energy distributions. Both corrected splitting methods are therefore effective for simulating Breit-Wheeler pair production across a wider range of photon energies, even with a relatively small number of photon macroparticles. For the rest of the paper, we focus on the Poisson approach, which is closer to the true physical picture at the cost of a higher runtime.
Figure 1: a) Rate of electron-positron pair production in a tightly focussed 1 PW beam against photon energy, comparing results without splitting (blue circles), with naive splitting (red ’x’s), with Poisson-sampling splitting (purple crosses), and with additive optical depths (green diamonds). b) Positron spectra predicted without splitting (blue) and with naive (red) and corrected splitting (purple), for photons at 400 MeV (i) and 1 GeV (ii).